OpenStudy (anonymous):

A jet plane, flying 110 mph faster than a propeller plane, travels 5280 miles in 4 hours less time than the propeller plane takes to fly the same distance. How fast does each plane fly?

OpenStudy (anonymous):

i hate word problems but i bet we can do it

OpenStudy (anonymous):

A jet plane, flying 110 mph faster than a propeller plane, travels 5280 miles in 4 hours less time than the propeller plane takes to fly the same distance. How fast does each plane fly?

OpenStudy (anonymous):

lets put \(x\) as the prop planes speed, so the speed of the jet is \(x+110\) since distance equals rate times time we know \[5280=xt\]for the prop plane and also \[5280=(x+110)(t-4)\] for the jet

OpenStudy (anonymous):

you can get rid of the \(t\) from the first equation and put \[t=\frac{5280}{x}\] put that in to the second one and get \[(x+110)(\frac{5280}{x}-4)=5280\]

OpenStudy (anonymous):

solve that happiness and find \(x=330\)

OpenStudy (anonymous):

thank you so much! how would I find the second speed?

OpenStudy (anonymous):

Nevermind I got it thanks!